RReturn to question 6 ! Part 4 of 15 Required information The Foundational 15 (LO2-1, LO2-2, LO2-3, LO2-4] (The following infoequired information The Foundational 15 [LO2-1, LO2-2, LO2-3, LO2-4] [The following information applies to the questions displayed below.] Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March): Molding Fabrication Total Estimated total machine-hours used 2,500 1,500 4,000 Estimated total fixed manufacturing overhead $ 10,750 $ 15,450 $ 26,200 Estimated variable manufacturing overhead per machine-hour $ 1.70 $ 2.50 Job P Job Q Direct materials $ 16,000 $ 9,500 Direct labor cost $ 23,400 $ 8,700 Actual machine-hours used: Molding 2,000 1,100 Fabrication 900 1,200 Total 2,900 2,300 Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month. Required: For questions 1-9, assume that Sweeten Company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments and Job P included 20 units and Job Q included 30 units. For questions 10-15, assume that the company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. Foundational 2-4 (Algo) 4. What was the total manufacturing cost assigned to Job P? (Do not round intermediate calculations.)
Answer #1

Molding overhead rate = (10750/2500)+1.7 = 6 per MH

Fabrication overhead rate = (15450/1500)+2.5 = 12.80 per MH

Job P
Direct material 16000
Direct labor 23400
Molding overhead (2000*6) 12000
Fabrication overhead (900*12.80) 11520
Total manufacturing cost 62920